# Binarized Simplicial Convolutional Neural Networks

### v1

@@ 4.1 Simplicial Convolution
where $\mathbf{U}_{k}$ is the eigenvector matrix and $\mathbf{\Lambda}_{k}$ is the eigenvalue matrix. The eigenvalues on the diagonals of $\mathbf{V}_{k}$ are arranged in increasing order and are referred to as frequencies and each eigenvector is the column of $\mathbf{U}_{k}$ appearing in the same order. The forward SFT is $\tilde{\mathbf{X}}_{k}=\mathbf{U}_{k}^{T}\mathbf{X}_{k}$, and the inverse SFT is $\mathbf{X}_{k}=\mathbf{U}_{k}\tilde{\mathbf{X}}_{k}$ respectively. A simplicial filter can be represented using a function on the frequencies matrix $h(\mathbf{\Lambda}_{k})$. The output $\boldsymbol{y}_{k}$ of a spectral simplicial convolution when the input is a one-dimensional simplicial vector $\boldsymbol{x}_{k}$:
$\boldsymbol{y}_{k}=\mathbf{U}_{k}h(\mathbf{\Lambda}_{k})\tilde{\mathbf{X}}_{k}=\mathbf{U}_{k}h(\mathbf{\Lambda}_{k})\mathbf{U}_{k}^{T}{\boldsymbol{x}}_{k}.$ (3)
The spectral convolution of a $k$-simplex with feature size $N_{k}$ relies on the eigendecomposition that has $O(N_{k}^{3})$ complexity and is numerically unstable when $N_{k}$ is large. Luckily, it has been shown that a series of polynomial summations approximate the above spectral approach, resulting in the spatial simplicial convolution [19]:
$\mathbf{U}_{k}h(\mathbf{\Lambda}_{k})\mathbf{U}_{k}^{T}\boldsymbol{x}_{k}=h(\mathbf{L}_{k})\approx\sum^{J}_{j=0}w_{k,j}\mathbf{L}^{j}_{k}\boldsymbol{x}_{k},$ (4)
@@
where $\mathbf{B}_{k}^{T}\boldsymbol{x}_{k-1}$ is the portion of $\boldsymbol{x}_{k}$ that is induced from $\boldsymbol{x}_{k-1}$, $\mathbf{B}_{k+1}\boldsymbol{x}_{k+1}$ is the portion of $\boldsymbol{x}_{k}$ that is induced from $\boldsymbol{x}_{k+1}$, and $\boldsymbol{x}_{k,h}$ is the harmonic component that cannot be induced [18]. The harmonic component $\boldsymbol{x}_{k,h}$ of $\boldsymbol{x}_{k}$ has the property of
@@
In other words, the Hodge decomposition essentially decomposes the Hodge Laplacian matrix $\mathbf{L}_{k}$ into the gradient, curl, and harmonic components [19].
In (4), we can assign specific filter weights to the lower and upper Hodge Laplacians because by definition, in (1), the Hodge Laplacian matrix can be split into the upper and the lower Laplacian. This leads to the (single feature) convolution operation seen in SCNN [21]:
@@
where $\gamma_{k,j}$ is the weights of the lower Hodge Laplacian and $\theta_{k,j}$ is the weights for the upper Hodge Laplacian. A term $\xi_{k}\mathbf{I}\boldsymbol{x}_{k}$ is added to the simplicial convolution because of the property shown in (6): the role of $\xi_{k}\mathbf{I}\boldsymbol{x}_{k}$ is to take into account of the harmonic component as it can only be processed independently from the upper or lower Laplacian [18]. Combining the simplicial convolution with nonlinearity, the forward propagation of an SCNN layer is
@@
where $\sigma()$ is the activation function, $p$ is the layer number, and $\boldsymbol{z}_{k}^{p}$ is the output of the $p^{th}$ simplicial layer [21]. In the SCNN, $\gamma_{k,j}$, $\theta_{k,j}$, and $\xi_{k}$ are the trainable weights to be optimized. Each $k$-simplex in a simplicial complex with order $K$ will have individually their own $k$-simplex SCNN.
@@
where $\Gamma_{k,j}$, $\Theta_{k,j}$, and $\Xi_{k,j}$ are the trainable weights with their sizes determined by the number of input and output features of layer $p$. The SCNN multi-variate version of the SCNN is also provided here as it was not given in the SCNN original literature [21]:
@@
4.2 Over-Smoothing Alleviation via Architectural Simplifications
First, notice that for $J>1$, we can intuitively acquire each summation term of length $J>1$ convolution by feeding our input into a series of concatenated $J$ length 1 convolution. In other words, one SCNN layer of length $J>1$ is equivalent to combining many SCNN layers of length $J=1$ without nonlinearity. To see this more clearly, let us consider the case of $J_{l}=J_{u}=2$ and break down (10):
@@ 4.2 Over-Smoothing Alleviation via Architectural Simplifications
$\mathbf{S}_{2}=\mathbf{Z}_{k,p}+\mathbf{L}_{k,l}^{2}\mathbf{Z}_{k,p}\mathbf{\Gamma}_{k,2}+\mathbf{L}_{k,u}^{2}\mathbf{Z}_{k,p}\mathbf{\Theta}_{k,1}.$ (13)
It is easy to see that $\mathbf{S}_{1}$ is just a length 1 convolution if we set $J=1$ in (9). Now, as with $\mathbf{S}_{2}$, it actually can be formulated as
$\mathbf{S}_{2}=\mathbf{L}_{k,l}\mathbf{Y}_{k,\hat{p}}\mathbf{\Gamma}_{k,2}+\mathbf{L}_{k,u}\mathbf{Y}_{k,\bar{p}}\mathbf{\Theta}_{k,2}+\mathbf{0}_{k}\mathbf{Z}_{k,p}$ (14)
@@
$\mathbf{Y}_{k,\hat{p}}=\mathbf{L}_{k,l}\mathbf{Z}_{k,p}\mathbf{I}_{k,1}+\mathbf{L}_{k,u}\mathbf{Z}_{k,p}\mathbf{0}_{k,1}+\mathbf{0}_{k}\mathbf{Z}_{k,p},$ (15)
@@
$\mathbf{Y}_{k,\bar{p}}=\mathbf{L}_{k,l}\mathbf{Z}_{k,p}\mathbf{0}_{k,1}+\mathbf{L}_{k,u}\mathbf{Z}_{k,p}\mathbf{I}_{k,1}+\mathbf{0}_{k}\mathbf{Z}_{k,p}.$ (16)
From the above equations, it is not difficult to check that for one SCNN layer with $J=2$, it is equivalent to having a combination of $4$ SCNN layers with $J=1$ stacked in two layers; a visualization of breaking down (11) is provided in Figure 1.
@@
In order to avoid this complexity of SCNN, our Bi-SCNN will be formed using purely length 1 simplicial convolution which gives us the advantage of alleviating over-smoothing. Let us look at the cause of over-smoothing. Each multiplication of the Hodge Laplacian is a 1-hop neighbor aggregation because, by definition, the non-diagonal elements of the Hodge Laplacian matrix are non-zero only when there is a neighbor in between $v_{i}$ and $v_{j}$ two elements $i$ and $j$. So, multiplying the Hodge Laplacian $\mathbf{L}_{k}$ with a feature vector $\boldsymbol{x}_{k}$ essentially aggregates all the neighborhood features that are 1-hop away, and doing it for $\mathbf{L}^{p}_{k}$ will aggregate from $p$-hops away. For the SCNN layer in (8), if $J$ is large enough, one single SCNN layer is possible to aggregate the entire simplex as each $j=1...J$ will aggregate from neighbors further and further. The over-smoothing problem is the exact analogy of what was seen in the GCNs [31] where the over-smoothed output has similar values among different elements in a simplex and causes the output to be indistinguishable. In the past, one simple yet effective solution to this problem on the graph neural network algorithms was to reduce the number of layers. In Bi-SCNN, we set the length of the filter to $J=1$, and treat the number of layers as a hyperparameter during training to reduce over-smoothing.
Figure 2: An illustration of a 3-layer Bi-SCNN network.

### v2

@@ 4.1 Simplicial Convolution and Topological Signal Processing
where $\mathbf{U}_{k}$ is the eigenvector matrix and $\mathbf{\Lambda}_{k}$ is the eigenvalue matrix. Following the graph Fourier transform convention, the eigenvalues are again assigned the notion of frequency. The eigenvalues on the diagonals of $\mathbf{V}_{k}$ are arranged in increasing order. Each eigenvector $\boldsymbol{u}_{i=1...N_{1}}$is the column of $\mathbf{U}_{k}$ and is arranged to appear in the same order as the eigenvalues. A few visual illustrations of the spectral components obtained from the SFT are shown in Figure 2.
The forward SFT projects a simplicial signal $\boldsymbol{x}_{k}$ into the spectral domain, thereby representing the signal in terms of frequency components:
$\tilde{\boldsymbol{x}}_{k}=\mathbf{U}_{k}^{T}\boldsymbol{x}_{k},$ (3)
where $\tilde{\boldsymbol{x}}_{k}$ is $\boldsymbol{x}_{k}$ transformed into the spectral domain. Similarly, the inverse SFT will invert the process of the forward SFT in which the signal $\boldsymbol{x}_{k}$ is being reconstructed from the spectral domain back to the spatial domain:
$\boldsymbol{x}_{k}=\mathbf{U}_{k}\tilde{\boldsymbol{x}}_{k}.$ (4)
(a) $\boldsymbol{u}_{1}$
(b) $\boldsymbol{u}_{20}$
(c) $\boldsymbol{u}_{N_{1}}$
Figure 2: Decomposing the edges of the graph with $N_{1}=38$ edges in Figure 1 using the SFT in (2) and $k=1$.
As we pointed out earlier, from the signal processing perspective, the modification of the signal will be conducted in the spectral domain, which requires the definition of a simplicial filter. A simplicial filter can be represented using a function on the frequencies matrix $h(\mathbf{\Lambda}_{k})$. Afterward, using the convolution property, applying this filter to the simplicial signal $\boldsymbol{x}_{k}$ can be achieved in the spectral domain by
$\boldsymbol{y}_{k}=\mathbf{U}_{k}h(\mathbf{\Lambda}_{k})\tilde{\boldsymbol{x}}_{k}=\mathbf{U}_{k}h(\mathbf{\Lambda}_{k})\mathbf{U}_{k}^{T}{\boldsymbol{x}}_{k},$ (5)
where $\boldsymbol{y}_{k}$ is the result of the simplicial convolution. For an input simplicial signal $\boldsymbol{x}_{k}$, different filters such as the low-pass filters or the high-pass filters can be applied using (5) to enhance desired features, suppress unwanted noise, or isolate specific components. Figure 3 provides an illustration of applying low-pass and high-pass filters to the edge signal $\boldsymbol{x}_{1}$.
(a) The edge signal in Figure 1 going through a low-pass filter.
(b) The edge signal in Figure 1 going through a high-pass filter.
Figure 3: An illustration of simplicial filtering on an edge signal using (5).
The spectral convolution of a $k$-simplex with feature size $N_{k}$ in (5) relies on the eigendecomposition that has $O(N_{k}^{3})$ complexity and is numerically unstable when $N_{k}$ is large. Luckily, it has been shown that a series of polynomial summations approximate the above spectral approach, resulting in the spatial simplicial convolution (Yang et al., 2022b, ):
$\mathbf{U}_{k}h(\mathbf{\Lambda}_{k})\mathbf{U}_{k}^{T}\boldsymbol{x}_{k}=h(\mathbf{L}_{k}){\boldsymbol{x}}_{k}\approx\sum^{J}_{j=0}w_{k,j}\mathbf{L}^{j}_{k}\boldsymbol{x}_{k},$ (6)
@@
where $\mathbf{B}_{k}^{T}\boldsymbol{x}_{k-1}$ is the portion of $\boldsymbol{x}_{k}$ that is induced from $\boldsymbol{x}_{k-1}$, $\mathbf{B}_{k+1}\boldsymbol{x}_{k+1}$ is the portion of $\boldsymbol{x}_{k}$ that is induced from $\boldsymbol{x}_{k+1}$, and $\boldsymbol{x}_{k,h}$ is the harmonic component that cannot be induced (Barbarossa and Sardellitti,, 2020). The harmonic component $\boldsymbol{x}_{k,h}$ of $\boldsymbol{x}_{k}$ has the property of
@@
In other words, the Hodge decomposition in (7) essentially decomposes the Hodge Laplacian matrix $\mathbf{L}_{k}$ into the gradient, curl, and harmonic components (Yang et al., 2022b, ).
4.2 Simplicial Convolution and TSP in Neural Networks
In (6), we can assign specific filter weights to the lower and upper Hodge Laplacians because by definition, in (1), the Hodge Laplacian matrix can be split into the upper and the lower Laplacian. This leads to the (single feature) convolution operation seen in SCNN (Yang et al., 2022a, ):
@@ 4.2 Simplicial Convolution and TSP in Neural Networks
where $\gamma_{k,j}$ is the weights of the lower Hodge Laplacian and $\theta_{k,j}$ is the weights for the upper Hodge Laplacian. A term $\xi_{k}\mathbf{I}\boldsymbol{x}_{k}$ is added to the simplicial convolution because of the property shown in (8): the role of $\xi_{k}\mathbf{I}\boldsymbol{x}_{k}$ is to take into account of the harmonic component as it can only be processed independently from the upper or lower Laplacian (Barbarossa and Sardellitti,, 2020). Combining the simplicial convolution with nonlinearity, the forward propagation of an SCNN layer is
@@
where $\sigma()$ is the activation function, $p$ is the layer number, and $\boldsymbol{z}_{k}^{p}$ is the output of the $p^{th}$ simplicial layer (Yang et al., 2022a, ). In the SCNN, $\gamma_{k,j}$, $\theta_{k,j}$, and $\xi_{k}$ are the trainable weights to be optimized. Each $k$-simplex in a simplicial complex with order $K$ will have individually their own $k$-simplex SCNN.
4.3 Simplicial Convolution Architectural Analysis
@@
where $\Gamma_{k,j}$, $\Theta_{k,j}$, and $\Xi_{k,j}$ are the trainable weights with their sizes determined by the number of input and output features of layer $p$. The SCNN multi-variate version of the SCNN is also provided here as it was not given in the SCNN original literature (Yang et al., 2022a, ):
@@
It is worth mentioning that the SCNN reduces to SNN (Ebli et al.,, 2020) if we set $\mathbf{\Xi}_{k}=0$ and $\mathbf{\Gamma}_{k,j}=\mathbf{\Theta}_{k,j}$. This essentially means that the SNN is the SCNN without Hodge decomposition (7). The multi-variate version of SNN is given as a reference:
$\mathbf{Z}_{k,p+1}=\sigma\left(\sum^{J_{l}}_{j=1}\mathbf{L}_{k}^{j}\mathbf{Z}_{k,p}\mathbf{\Gamma}_{k,j}\right).$ (13)
For simplicity of analysis, let us assume the length of simplicial convolutions in (13) is $J_{u}=J_{l}=J$. Notice that for $J>1$, we can intuitively acquire each summation term of length $J>1$ convolution by feeding our input into a series of concatenated $J$ length 1 convolution. In other words, one SCNN layer of length $J>1$ can be represented as the combination of multiple SCNN layers of length $J=1$ without nonlinearity. To see this more clearly, let us consider the case of $J_{l}=J_{u}=2$ and break down (13):
@@
$\mathbf{S}_{2}=\mathbf{Z}_{k,p}+\mathbf{L}_{k,l}^{2}\mathbf{Z}_{k,p}\mathbf{\Gamma}_{k,2}+\mathbf{L}_{k,u}^{2}\mathbf{Z}_{k,p}\mathbf{\Theta}_{k,2}.$ (16)
It is easy to see that $\mathbf{S}_{1}$ is just a length 1 convolution, which is equivalent to the case when we set $J=1$ in (11). Now, as for $\mathbf{S}_{2}$ in (16), it actually can be expressed as
$\mathbf{S}_{2}=\mathbf{L}_{k,l}\hat{\mathbf{Y}}_{k,p}\mathbf{\Gamma}_{k,2}+\mathbf{L}_{k,u}\tilde{\mathbf{Y}}_{k,p}\mathbf{\Theta}_{k,2}+0\mathbf{Z}_{k,p}$ (17)
@@
$\hat{\mathbf{Y}}_{k,p}=\mathbf{L}_{k,l}\mathbf{Z}_{k,p}\mathbf{I}_{k}+\mathbf{L}_{k,u}\mathbf{Z}_{k,p}0+0\mathbf{Z}_{k,p},$ (18)
@@
$\tilde{\mathbf{Y}}_{k,p}=\mathbf{L}_{k,l}\mathbf{Z}_{k,p}0+\mathbf{L}_{k,u}\mathbf{Z}_{k,p}\mathbf{I}_{k}+0\mathbf{Z}_{k,p}.$ (19)
@@
From the above equations, it is not difficult to check that for one SCNN layer with $J=2$, it is equivalent to having a combination of $4$ SCNN layers with $J=1$ stacked in two layers. The reason behind this is that when $J>1$, for each $j$ that is greater than $1$, the simplicial length $J$ convolution in (14) will introduce a calculation of either $\mathbf{L}^{j}_{k,l}$ or $\mathbf{L}^{j}_{k,u}$. From the matrix multiplication perspective, using the lower adjacency aggregation $\mathbf{L}^{j}_{k,l}\mathbf{Z}_{k,p}=\mathbf{L}^{j-1}_{k,l}(\mathbf{L}_{k,l})\mathbf{Z}_{k,p}$ as an example, we can see that for $J=2$, it requires (14) to do the simplicial length 1 convolution twice to calculate the results. A similar argument for obtaining the results for upper adjacency can be made. As a result, setting $J=2$ in (14) is essentially doing the computation of $4$ (14) with $J=1$, which introduces high computational complexity and over-smoothing when the length of convolution $J$ or the number of layers $P$ is high. A visualization of the above breaking down of (14) is provided in Figure 4 for the case of $J=2$. Intuitively, for an SNN with a generic length $J$ that is large, a length $J$ SCNN layer can be decomposed into many length 1 SCNN layers using the above analysis. In other words, when $J$ is large, one layer of SCNN intuitively incorporates many length 1 SCNNs, which leads to the high model and computational complexity of SCNN.
@@
In addition to the complexity issue pointed out above, the SCNN is also prone to over-smoothing. To see this, let us revisit the SCNN in (13). Each multiplication of the Hodge Laplacian in SCNN is a 1-hop neighbor aggregation because, by definition, the non-diagonal elements of the Hodge Laplacian matrix are non-zero only when there is a neighbor in between $v_{i}$ and $v_{j}$ two elements $i$ and $j$. So, multiplying the Hodge Laplacian $\mathbf{L}_{k}$ with a feature vector $\boldsymbol{x}_{k}$ essentially aggregates all the neighborhood features that are 1-hop away, and doing it for $\mathbf{L}^{j}_{k}$ will aggregate from $j$-hops away. For the SCNN layer in (10), if $J$ is large, one single SCNN layer is possible to aggregate the entire simplex as each $j=1...J$ will aggregate from neighbors further and further, which makes SCNN prone to over-smoothing of the features. The over-smoothing problem is the exact analogy of what was seen in the GCNs (Liu et al.,, 2020) where the over-smoothed output has similar values among different elements in a simplex and causes the output to be indistinguishable.
Before we proceed with the Bi-SCNN, let us briefly discuss the Simplicial MPNN. In short, the Simplicial MPNN increased the representation ability of SCNN by individually defining the aggregation of simplex using the concept of message. There are 4 types of Message Passing on a simplicial complex: the upper adjacency, the lower adjacency, the boundary adjacency, and the coboundary adjacency (Bodnar et al.,, 2021). In Simplicial MPNN, each of the adjacencies can be modeled as trainable weights, and the weights can be learned and updated through backpropagation. As so, the number of trainable weights in a Simplicial MPNN layer is proportional to the number of simplices in the particular simplicial order of the adjacency of choice. The significant increase in the number allows the Simplicial MPNN to have better representation power compared to GCN, SNN, and SCNN when properly trained. The SNN in (12) and SCNN in (13) can be obtained by specifying MPNN using upper and lower adjacencies. Here is a realization of the SCNN using Simplicial MPNN:
$\mathbf{{Z}}_{k,{p+1}}=\sigma\left(\mathbf{B}_{k}^{T}\text{diag}(\boldsymbol{\gamma}_{k,p})\mathbf{B}_{k}\mathbf{{Z}}_{k,{p}}+\mathbf{B}_{k+1}\text{diag}(\boldsymbol{\theta}_{k,p})\mathbf{B}_{k+1}^{T}\mathbf{{Z}}_{k,{p}}\right),$ (20)
where $\boldsymbol{\gamma}_{k,p}$ and $\boldsymbol{\theta}_{k,p}$ are the trainable parameters Even though MPNN has high representation power, it comes at a cost of significantly increased parameter size, causing the runtime and computational complexity for training to be high. This also means that the Simplicial MPNN shares the same drawbacks as SNN and SCNN we pointed out earlier.
4.4 Simplicial Feature Binarization
Let us now examine how to simultaneously reduce the model complexity, enhance the time efficiency, and maintain the effectiveness of its simplicial embedding of the simplicial-convolution-based architectures. We will address each of the identified challenges systematically.
First, the complexity of the SCNN causes slow computation and over-smoothing. In the past, one simple yet effective solution to the over-smoothing problem of graph neural network algorithms was to reduce the number of layers. In order to avoid the complexity of SCNN, the proposed Bi-SCNN will be formed using purely length 1 simplicial convolution. In Bi-SCNN, we set the length of the filter to $J=1$, and treat the number of layers as a hyperparameter during training to reduce over-smoothing. It has been demonstrated on GNNS with node-based graph data that similar reductions in computation can enhance the scalability of neural networks and help mitigate over-smoothing (Kipf and Welling,, 2016).
To reduce the model complexity, other than simply reducing the number of parameters, one of the many solutions is Binarization. One example of binarization is the Bi-GCN, where both the trainable weights and the features are binarized; the forward propagation is done by binary operations bit count and XOR to combine the binarized weights and features into a new node embedding. Aside from the fact that these node-based algorithms are not suitable to be deployed on simplicial complexes, there are a few additional problems to be addressed. The Bi-GCN binarizes both the weights and features, leading to reduced performance under certain circumstances. Also, even though Bi-GCN uses binary operations to optimize the computation cost, the forward propagation is indifferent and mathematically identical compared with the original GCN. Another approach is seen in the Graph-Sign algorithms (Yan et al.,, 2022), which delivers a sign-error update at each iteration that has a low run time and a low complexity. The drawback of the Graph-Sign algorithm is that it operates on a predefined bandlimited filter, which may require prior expert knowledge to obtain.
In Bi-SCNN, we propose a new architecture where we combine feature normalization and feature binarization together with simplicial filters; the nonlinearity is achieved via feature normalization and feature binarization, allowing forward and backward propagation to be efficient and accurate. Let us revisit the multi-variate simplicial convolution in (13). Take an approach that resembles what was seen in the Bi-GCN (Wang et al., 2021b, ) by following the graph-based feature binarization step in Bi-GCN, we define the feature binarization of simplicial data $\mathbf{X}_{k}$ on $k$-simplex $\mathcal{S}_{k}$ as
where $\circ$ is the Hadamard product, Sign()()( ) is the Sign function and $\mathbf{\bar{X}_{k}}$ is the (weighted) binarized simplicial features. As a reference, the Sign() function is defined as
[…]
The matrix $\mathbf{M}_{k}$ acts as the weights for the binarized features and is formed by a feature normalization:
${m_{i}}=\frac{norm_{1}(\mathbf{r}_{k})}{d_{k}},$ (23)
where $norm_{1}(\mathbf{r}_{k})$ is the $l_{1}$-norm done on the $i^{th}$ row of $\mathbf{X}_{k}$ ($\mathbf{r}_{k}$ in (23)) and $d_{k}$ is the number of features in $\mathcal{S}_{k}$. All elements of the same row in $\mathbf{M}_{k}$ have the same value, but this is only for mathematical representation convenience. In practice, $\mathbf{M}_{k}$ is stored as a vector $\boldsymbol{m}_{k}$ and is broadcasted using element-wise multiplications when used to multiply with $\text{Sign}(\mathbf{X}_{k})$. Notice that the Sign() function only outputs binary values $-1$ and $+1$ assuming that we ignore the zeros, this means that the size of the binarized features $\mathbf{\bar{X}_{k}}$ is reduced in practice when compared to $\mathbf{X}_{k}$ (Wang et al., 2021b, ).
$\mathbf{\bar{Z}}_{k,{p+1}}=\sigma\left(\mathbf{L}_{k,l}\mathbf{\bar{Z}}_{k,{p}}\mathbf{\Gamma}_{k,{p}}+\mathbf{L}_{k,u}\mathbf{\bar{Z}}_{k,{p}}\mathbf{\Theta}_{k,{p}}+\mathbf{\bar{Z}}_{k,{p}}\mathbf{\Xi}_{k,{p}}\right).$ (24)
Looking at the expression in (24), aside from benefiting from the model size reduction, (24) still follows the exact same forward propagation strategy as SCNN. Under unfavorable scenarios, (24) could potentially only offer a marginal speed advantage, resulting in approximately equivalent execution times. We would like to improve the run speed of this Binary setup further by propagating only the Sign() part and further reducing the computation. The motivation behind this is that the Sign only has 3 outputs, with the output 0 rarely appearing in practice, so propagating only 1s and -1s significantly reduces the computation. Using the two properties of Hadamard products $\mathbf{A}\circ\mathbf{B}=\mathbf{B}\circ\mathbf{A}$ and $\mathbf{A}\circ\mathbf{(B+C)}=\mathbf{A}\circ\mathbf{B}+\mathbf{A}\circ\mathbf{C}$, with some algebraic derivations, equation (24) becomes
$\begin{split}\mathbf{\bar{Z}}_{k,{p+1}}=&\mathbf{M}_{k,p}\circ(\mathbf{L}_{k,l}\text{Sign}(\mathbf{\bar{Z}}_{k,{p}})\mathbf{\Gamma}_{k,{p}}+\\ &\mathbf{L}_{k,u}\text{Sign}(\mathbf{\bar{Z}}_{k,{p}})\mathbf{\Theta}_{k,{p}}+\text{Sign}(\mathbf{\bar{Z}}_{k,{p}})\mathbf{\Xi}_{k,{p}}).\end{split}$ (25)
Notice that the $l_{1}$-norm is nonnegative, so at each layer, $\mathbf{M}_{k,p}$ can be viewed as $\mathbf{M}_{k,p}=$ ReLu($\mathbf{M}_{k,p}$). The input of the $p+1^{th}$ layer in (25) will binarize the output of the $p^{th}$ layer regardless of the magnitude, so we can safely input only the binarized simplicial features (without the weights $\mathbf{M}_{k}$) to the $p+1^{th}$ layer when $p>1$. Also, we should point out that the Sign() function is non-linear. In this way, the nonlinearity in both the Sign() function and the $l_{1}$-norm can serve as activations in forward propagation. Additionally, both paths will benefit from a reduction in computational complexity, which we will be analyzing in the next subsection.